# Unveiling galaxy chemical enrichment mechanisms out to cosmic dawn from direct determination of O & Ar abundances from JWST/NIRSPEC spectroscopy

### v1

@@ Abstract
Galaxy chemical enrichment mechanisms have primarily been constrained by [/Fe] and [Fe/H] measurements of individual stars and integrated light from stellar populations. However such measurements are limited at higher redshifts (z). Recently, we proposed an analogous diagram of the oxygen-to-argon abundance ratio, log(O/Ar), vs Ar abundance, 12+log(Ar/H), as a new diagnostic window for emission nebulae. In this Letter, using robust line flux measurements including temperature sensitive auroral lines, we present direct determination of O and Ar abundances in nine SFGs from JWST/NIRSPEC spectra at z1.3–7.7, and two more with Keck/MOSFIRE spectra at z2.2. Utilising their positions on the log(O/Ar) vs 12+log(Ar/H) plane, we present the first inference of galaxy chemical enrichment mechanisms from an ensemble of galaxies. The SFGs at z1.3–3.4 are consistent with the solar neighbourhood galactic chemical enrichment models of the Milky Way Galaxy that are driven by core-collapse and Type Ia supernovae. Such enrichment mechanisms thus occur at least out to z. However, the highest-redshift SFGs (z3.6–7.7) have very low log(O/Ar) values, revealing a different enrichment process at z. Such low log(O/Ar) values may be caused by a rapid but intermittent star-formation and/or additional sources. The new diagnostic window for SFGs enables us to reveal the unique fingerprints of galaxy chemical enrichment out to cosmic dawn.
@@ 1 Introduction
Since the break of cosmic dawn, the interstellar medium (ISM) of galaxies has been continually enriched by the birth and death of stars. The bulk of our understanding of galaxy chemical enrichment (Tinsley, 1980; Pagel, 1997; Kobayashi et al., 2020a; Matteucci, 2021) stems from spectroscopic observations of individual stars in our Milky Way (MW) and studies of nearby galaxies. Deep absorption line spectra of MW stars enabled determination of their [Fe/H] and [/Fe], revealing the chemical composition of the ISM at the time of their birth (e.g. Edvardsson et al. 1993; Fuhrmann 1998; Hayden et al. 2015). The stars showing the highest [/Fe] values are thought to have formed at the earliest times from ISM that had only been enriched by core-collapse (CCSNe). Once Type Ia supernovae (SNe Ia) explosions begin, more Fe is released to the ISM than previously, causing a decreasing trend in [/Fe] vs [Fe/H] for the subsequent generations of stars.
@@
Recently, the log(O/Ar) vs 12 + log(Ar/H) plane for emission nebulae was found to be analogous to the [/Fe] vs [Fe/H] plane for stars (Arnaboldi et al., 2022). This was based on the analysis of emission-line spectra of planetary nebulae and HII regions surveyed in the Andromeda galaxy (M 31; Bhattacharya et al. 2019, 2021) where temperature sensitive auroral lines had been observed (Bhattacharya et al., 2022; Esteban et al., 2020). Like Fe, SNe Ia also preferentially produce more Ar than light -elements like O, whereas CCSNe produce near-constant log(O/Ar), see Kobayashi et al. (2020a). Based on this concept, Arnaboldi et al. (2022) introduced the log(O/Ar) vs 12 + log(Ar/H) plane using planetary nebulae to reveal the chemical enrichment history of M 31 with high star-formation at early times ( Gyr ago) and gas-infall 2–4 Gyr ago. Kobayashi et al. (2023) then showed that with well-constrained GCE models, the log(O/Ar) vs 12 + log(Ar/H) plane can be connected to the [/Fe]-[Fe/H] plane.
@@
These elemental abundances trace the state of the cumulative chemical enrichment of its ISM by previous generations of stars. As we move to higher redshifts, the chemical abundances in galaxies increasingly map the chemical enrichment from the very early generations of stars. As the relative contribution of CCSNe and SNeIa varies for the production of different elements (see Kobayashi et al., 2020a), we may use the relative abundances of different elements to decipher the state of chemical enrichment of each SFG at their observed redshift.
Rogers et al. (2024) and Welch et al. (2024) determined log(O/Ar) values for two SFGs at z1.3 and z3 respectively. Rogers et al. (2024) utilised the aforementioned results (Kobayashi et al., 2020a; Arnaboldi et al., 2022) to interpret the state of chemical enrichment of their individual SFG with a super-solar log(O/Ar) value as being primarily enriched by CCSNe. Individual SFGs may occupy different positions in the log(O/Ar) vs 12 + log(Ar/H) plane, and thereby exhibit different states of chemical enrichment. However, by determining the positions of an ensemble of high-z SFGs in this plane, we can constrain the mechanisms that drive early galaxy chemical enrichment.
@@
Name | z | c(H) | Line Flux (S/N) | Source | Reference | |||
[OIII] 4363 Å | [ArIV] 4711 Å | [ArIV] 4740 Å | [ArIII] 7136 Å | 1D spectra | Line flux | |||
ERO 10612 | 7.66 | (8.3) | (3.27) | (3.33) | – | MASTb | Curti et al. (2023) | |
CEERS 1536 | 5.038 | (4.24) | – | – | (3.96) | CEERS DR07c | Nakajima et al. (2023) | |
GLASS 150029a | 4.584 | (6.64) | – | – | (3.89) | MASTb | Nakajima et al. (2023) | |
CEERS 1665a | 4.488 | (3.3) | – | – | (8.54) | CEERS DR07c | Sanders et al. (2024) | |
CEERS 1651 | 4.382 | (3.77) | – | (8.94) | (5.32) | CEERS DR07c | Sanders et al. (2024) | |
GLASS 40066 | 4.02 | (4.71) | – | – | (4.95) | MASTb | Isobe et al. (2023) | |
JADES 19519 | 3.604 | (3.8) | – | – | (3.22) | Bunker et al. (2023) | Morishita et al. (2024) | |
CEERS 11088 | 3.302 | (3.45) | – | – | (6.55) | CEERS DR07c | Sanders et al. (2024) | |
CEERS 3788 | 2.295 | (4.78) | – | – | (4.94) | CEERS DR07c | Sanders et al. (2024) |
@@ 2 Data and abundance determination
2.1 Galaxy sample
We build a sample of high-redshift (z1) galaxies with O and Ar abundances determined directly via the detection of temperature sensitive auroral lines. For this purpose, we identify galaxies that have O abundances published already in the literature from their [OIII] 4363 Å line detected in JWST/NIRSPEC observations.
We first searched for their publicly available flux and wavelength calibrated 1D JWST/NIRSPEC grating spectra and verified if their [OIII] 4363 Å line was detected with with signal-to-noise S/N in these spectra. We then check if ionised Ar lines were also present (with S/N). This procedure resulted in a sample of nine such galaxies for which we obtain direct determination of both O and Ar abundances. The flux determination procedure for these sources is laid out in Section 2.2. The [OIII] 4363 Å and ionised Ar line fluxes are listed in Table 1 for these nine galaxies, along with the literature sources where their [OIII] 4363 Å flux and O abundance were first published and where other relevant line flux measurements are also available. The reliability of the flux calibration and its effect on the determined O and Ar abundances are discussed further in Appendix A. Two of these galaxies (GLASS 150029 and CEERS 1665), having relatively broad H lines, have been reported as AGN hosts111As the AGN is unresolved for the two sources, it remains unclear if it is responsible for all their emission-line fluxes. We thus include them in our abundance determination analysis assuming a star-forming ionising source but conservatively they are not considered for the ensuing interpretation. by Harikane et al. (2023) while the other seven are SFGs.
@@ 2.1 Galaxy sample
We further include four additional high-redshift SFGs in our sample whose relevant line flux measurements, including ionised O and Ar lines, are published in the literature, though their spectra were not available publicly. Two of these SFGs had their O and Ar abundances determined directly based on their [OIII] 4363 Å line detection (Rogers et al., 2024; Welch et al., 2024). For two SFGs at z2.2, their direct determination of O abundances were based on Keck/MOSFIRE spectra but with temperature sensitive [OII] 7322,7332 Å line detections, where [ArIII] 7136 Å line fluxes were also reported (Sanders et al., 2023). We use the published line-flux measurements for these four additional SFGs to determine their O & Ar abundances.
2.2 Line-flux measurement
For the galaxies where we had publicly available flux and wavelength calibrated 1D JWST/NIRSPEC grating spectra, the fluxes for the relevant emission-lines were measured using the automated line fitting algorithm, ALFA (Wesson, 2016), using their published redshift as an input parameter. After subtracting a globally fitted continuum, ALFA derives fluxes by optimizing the parameters of Gaussian fits to the line profiles using a genetic algorithm; line-blending are also taken into account (see Appendix B). The publicly available spectra of the nine galaxies showing the [OIII] 4363 Å lines and the ionised Ar lines (either [ArIII] 7136 Å and/or[ArIV] 4711,4740 Å) are shown in Figure 1. Note that for ERO 10612 (), the [ArIII] 7136 AA line lies beyond the red wavelength limit of NIRSPEC.
We used stacking techniques to increase the S/N of the observed spectra confirming the identification of the [OIII] 4363 Å and [ArIII] 7136 lines (Arnaboldi et al., 2002). Appendix A provides further detailed information. For each source, the flux ratios of these lines with respect to H, along with their S/N, are noted in Table 1. All other lines of interest, not listed in Table 1, are significantly brighter than [OIII] 4363 Å line and have been published previously (with almost no measured difference) in the references listed in the same table.

### v2

# Unveiling galaxy chemical enrichment mechanisms out to z8 from direct determination of O & Ar abundances from JWST/NIRSPEC spectroscopy

@@ Abstract
Galaxy chemical enrichment mechanisms have primarily been constrained by [/Fe] and [Fe/H] measurements of individual stars and integrated light from stellar populations. However such measurements are limited at higher redshifts (z). Recently, we proposed an analogous diagram of the oxygen-to-argon abundance ratio, log(O/Ar), vs Ar abundance, 12+log(Ar/H), as a new diagnostic window for emission nebulae. In this Letter, using line flux measurements including temperature sensitive auroral lines, we present direct determination of O and Ar abundances in nine star-forming galaxies (SFGs) from JWST/NIRSPEC spectra at z1.3–7.7, and two more with Keck/MOSFIRE spectra at z2.2. Utilising their positions on the log(O/Ar) vs 12+log(Ar/H) plane, we present the first inference of galaxy chemical enrichment mechanisms from an ensemble of galaxies. Seven SFGs at z1.3–4 are consistent with the Milky Way solar neighbourhood galactic chemical enrichment models that are driven by core-collapse and Type Ia supernovae in a self-regulated manner. Such enrichment mechanisms thus occur at least out to z. However, four higher-redshift SFGs (z3.6–7.7) have lower log(O/Ar) values, revealing potentially different enrichment paths becoming important at z. Such log(O/Ar) values may be caused by physical mechanisms such as a rapid but intermittent star-formation and/or additional enrichment sources. This new diagnostic window for SFGs enables us to reveal the unique fingerprints of galaxy chemical enrichment out to cosmic dawn.
@@ 1 Introduction
Since the break of cosmic dawn, the interstellar medium (ISM) of galaxies has been continually enriched by the birth and death of stars. The bulk of our understanding of galaxy chemical enrichment (Tinsley, 1980; Pagel, 1997; Kobayashi et al., 2020a; Matteucci, 2021) stems from spectroscopic observations of individual stars in our Milky Way (MW) and studies of nearby galaxies. Deep absorption line spectra of MW stars enabled determination of their [Fe/H] and [/Fe], revealing the chemical composition of the ISM at the time of their birth (e.g. Edvardsson et al. 1993; Fuhrmann 1998; Hayden et al. 2015; Imig et al. 2023). The stars showing the highest [/Fe] values are thought to have formed at the earliest times from ISM that had only been enriched by core-collapse (CCSNe). Once Type Ia supernovae (SNe Ia) explosions begin, more Fe is released to the ISM than previously, causing a decreasing trend in [/Fe] vs [Fe/H] for the subsequent generations of stars.
@@
Recently, the log(O/Ar) vs 12 + log(Ar/H) plane for emission nebulae was found to be analogous to the [/Fe] vs [Fe/H] plane for stars (Arnaboldi et al., 2022). This was based on the analysis of emission-line spectra of planetary nebulae and HII regions surveyed in the Andromeda galaxy (M 31; Bhattacharya et al. 2019a, b, 2021, 2023) where temperature sensitive auroral lines had been observed (Bhattacharya et al., 2022; Esteban et al., 2020). Like Fe, SNe Ia also preferentially produce more Ar than light -elements like O, whereas CCSNe produce near-constant log(O/Ar), see Kobayashi et al. (2020a). Based on this concept, Arnaboldi et al. (2022) introduced the log(O/Ar) vs 12 + log(Ar/H) plane using planetary nebulae to reveal the chemical enrichment history of M 31 with high star-formation at early times ( Gyr ago) and gas-infall 2–4 Gyr ago. Kobayashi et al. (2023) then showed that with well-constrained GCE models, the log(O/Ar) vs 12 + log(Ar/H) plane can be connected to the [/Fe]-[Fe/H] plane.
@@
These elemental abundances trace the state of the cumulative chemical enrichment of its ISM by previous generations of stars. As we move to higher redshifts, the chemical abundances in galaxies increasingly map the chemical enrichment from the very early generations of stars. As the relative contribution of CCSNe and SNe Ia varies for the production of distinct elements (see Kobayashi et al., 2020a), we may use the relative abundances of such elements to decipher the state of chemical enrichment of each SFG at their observed redshift.
Rogers et al. (2024) and Welch et al. (2024) determined log(O/Ar) values for two SFGs at z3 and z1.3 respectively. Rogers et al. (2024) utilised the aforementioned results (Kobayashi et al., 2020a; Arnaboldi et al., 2022) to interpret the state of chemical enrichment of their individual SFG with a super-solar log(O/Ar) value as being primarily enriched by CCSNe. Individual SFGs may occupy different positions in the log(O/Ar) vs 12 + log(Ar/H) plane, and thereby exhibit different states of chemical enrichment. However, by determining the positions of an ensemble of high-z SFGs in this plane, we can constrain the mechanisms that drive early galaxy chemical enrichment.
@@
| Name | [OII] | [OII] | H | H | [OIII] | [ArIV] | [ArIV] | [OIII] | [OIII] | H | [SII] | [SII] | [ArIII] |
| 3726 Å | 3729 Å | 4102 Å | 4340 Å | 4363 Å | 4711 Å | 4740 Å | 4959 Å | 5007 Å | 6563 Å | 6717 Å | 6731 Å | 7136 Å | |
| ERO 10612a | 14.4 3.9 | – | 20.9 3.6 | 50.8 7.4 | 24.8 3.7 | 14.1 4.1 | 20.3 5.1 | 274.7 9.3 | 869.3 14.9 | – | – | – | – |
| CEERS 1536a | 63.8 12.2 | – | – | 53.7 12.1 | 32.2 9.1 | – | – | 252.8 14.3 | 795.2 27.3 | 628.1 37.8 | – | – | 31.4 9.9 |
| GLASS 150029 | 17.8 4.1 | 20.7 4.0 | 25.5 3.2 | 39.6 4.6 | 18.4 3.5 | – | – | 232.8 9.1 | 658.8 15.2 | 464.3 11.9 | – | – | 16.6 3.2 |
| CEERS 1665 | 51.2 2.6 | 35.2 2.6 | 11.8 1.7 | 38.3 3.4 | 6.7 2.1 | – | – | 233.8 7.9 | 733.6 9.3 | 518.6 8.9 | 31.0 3.1 | 20.3 3.9 | 14.8 3.9 |
| CEERS 1651b | – | – | – | 26.8 5.9 | 11.4 3.6 | – | – | 217.3 8.9 | 636.6 9.2 | – | – | – | 28.4 9.2 |
| GLASS 40066c | – | – | 20.1 1.7 | 36.8 2.8 | 7.8 1.3 | – | – | 251.9 5.5 | 768.8 13.4 | 409.5 6.2 | 20.2 4.2 | – | 13.3 2.2 |
| JADES 19519 | 37.9 6.1 | 53.3 5.9 | – | 26.2 4.1 | 20.1 4.0 | – | – | 227.9 10.5 | 678.2 9.2 | 511.2 45.1 | 39.6 8.3 | – | 16.7 3.9 |
| CEERS 11088 | 52.0 3.1 | 47.0 3.4 | 14.6 3.9 | 38.0 3.2 | 8.8 2.7 | – | – | 193.6 3.9 | 623.3 7.6 | 696.6 22.7 | 58.3 7.1 | 25.7 7.0 | 22.9 5.0 |
| Q2343-D40c | – | – | – | 39.7 2.9 | 4.0 1.6 | – | – | 225.4 2.3 | 707.5 6.9 | 555.7 37.8 | 32.1 2.9 | 20.4 2.9 | 15.8 2.1 |
| CEERS 3788 | 41.7 1.9 | 45.9 1.9 | 16.0 2.8 | 32.2 2.5 | 8.8 1.6 | – | – | 237.0 5.4 | 701.0 8.2 | 608.7 9.5 | 37.2 5.6 | 23.3 5.8 | 19.8 3.5 |
@@ 2 Data and abundance determination
2.1 Emission-line galaxy sample from MAST
We built a sample of high-redshift (z1) galaxies with O and Ar abundances determined directly via the detection of temperature sensitive auroral lines. For this purpose, we first identified galaxies that have O abundances published already in the literature from their [OIII] 4363 Å line detected in JWST/NIRSPEC MOS observations. We then searched for their publicly available111From the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. 1D JWST/NIRSPEC grating spectra, and verified, via the publicly available automated line fitting algorithm ALFA software (Wesson, 2016), whether their [OIII] 4363 Å line was detected with signal-to-noise S/N in these spectra, using their previously published redshift as an input parameter. We also checked with ALFA whether the ionised Ar lines (either [ArIII] 7136 Å or [ArIV] 4711,4740 Å) were detected with S/N in these spectra. For the reliability of the flux and wavelength calibration, please see further in Appendix A.
This procedure resulted in a sample of ten galaxies for which all the emission lines required for direct abundance determination are present with S/N, see Table 1. The selected spectra cover a wide range of exposure times (0.85 – 29.2 hrs) depending on the parent survey; they are shown in Figure 1a. We note that the spectra of GLASS 150029 and GLASS 40066 have higher spectral resolution than those of the other galaxies, as their observations were carried out with the higher resolution G235H and G395H gratings, while the other galaxies were observed with the medium resolution G235M and G395M gratings. For the galaxy ERO 10612, two sets of observations were available from the MAST archive, acquired at different telescope rotation angles. For our analysis, we used the exposure with strong emission lines in the spectra that have symmetric line profiles while the other one with asymmetric line profiles is not used. All the utilized spectra in this study can be accessed via https://url.uk.m.mimecastprotect.com/s/683ACDYyMuy6PYXcWflTjmen3?domain=dx.doi.org (catalog DOI:10.17909/yj93-nc36) from MAST. Two of these galaxies at z4.5 (GLASS 150029 and CEERS 1665), having relatively broad H lines, have been reported as candidate AGN hosts222As the AGN is unresolved for the two sources, it remains unclear if it is responsible for all their emission-line fluxes. We thus include them in our abundance determination analysis assuming a star-forming ionising source but conservatively they are not considered for the ensuing interpretation. by Harikane et al. (2023) while the other eight are SFGs.
ALFA measures emission line fluxes from galaxy spectra, after subtracting a globally fitted continuum, by optimizing the parameters of Gaussian fits to the line profiles using a genetic algorithm; line-blending are also taken into account (see Appendix B). For faint lines, the wavelength positions and line profile shape, constrained from stronger emission lines, are used as additional discriminants from noise peaks reaching comparable counts/fluxes. As an example, Figure 2a shows the 1D spectrum of CEERS 1651 around its [ArIII] 7136 Å line. The ionised Ar line as well as the He I 7281 Å line (while not of interest in this work) have gaussian-like profiles and are detected as lines by ALFA. No line is detected at the position of the He I 7065 Å line (also not of interest in this work) as the S/N at this position is less than 3. Noise peaks showing non-gaussian line profiles are thus rejected by ALFA, even if the S/N at their positions is greater than 3.
Nevertheless, we will further verify both [OIII] 4363 Å and ionised Ar line detections by inspecting their 2D spectra as well as via a stacking analysis for the galaxies with faint [ArIII] 7136 Å lines, see Section 2.2.
2.2 Emission line verification
In Figure 1a, we show the 2D traces of the spectra for the 10 galaxies selected from the MAST archive, at the rest wavelength of the [OIII] 4363 Å lines and ionised Ar lines. The [OIII] 4363 Å lines are clearly visible in the 2D traces with varying intensity for all galaxies, with the faintest emission seen for CEERS 11088.
The ionised Ar lines are also visible in the 2D traces for all galaxies. We note that for ERO 10612, GLASS 150029 and CEERS 1651 there are noise peaks of similar intensity in the 2D traces. Along with CEERS 1536 & CEERS 1665, these constitute the highest redshift () galaxies in our sample (see Table 2). For these five galaxies, we additionally investigate the 2D distribution of their flux to flux error ratio, Flux/, around the rest-frame wavelengths of their detected ionised Ar lines, see Figure 1b.
For each galaxy, the Flux/ 2D traces display peaks at the wavelengths of the ionised Ar lines ([ArIV] 4711,4740 Å for ERO 10612 and [ArIII] 7136 Å for the others) within the region bracketed by the red lines in Figure 1b, where the 1D spectrum is extracted. Some off-centered peaks are also visible for CEERS 1536 and CEERS 1665 at Å bluer wavelength than the detected Ar line in Figure 1b. These peaks have sharp profiles which are different from that of the faint line emissions, which are spread on several pixels instead, and better centered within the extraction region.
| Name | z | log(M∗) | log(sSFR) | tSF | c(H) | Te | ne | 12+log(O/H) | 12+log(Ar/H) | log(O/Ar) |
| [M⊙] | [yr-1] | [Myr] | [K] | [cm-3] | ||||||
| ERO 10612 | 7.66 | |||||||||
| CEERS 1536 | 5.038 | 1000c | ||||||||
| GLASS 150029a | 4.584 | |||||||||
| CEERS 1665a | 4.488 | |||||||||
| CEERS 1651 | 4.382 | 1000c | ||||||||
| GLASS 40066 | 4.02 | 1000c | ||||||||
| JADES 19519 | 3.604 | |||||||||
| CEERS 11088 | 3.302 | |||||||||
| Q2343-D40 | 2.963 | – | – | – | ||||||
| CEERS 3788 | 2.295 | |||||||||
| COSMOS 19985b | 2.188 | |||||||||
| COSMOS 20062b | 2.185 | |||||||||
| SGAS1723+34b | 1.329 |
Note. — aFor these galaxies, the [OII] 3726, 3729 Å doublet appears blended and its total flux noted in the [OII] 3726 Å column.
bFor this galaxy, the [OII] 3726, 3729 Å doublet, H and [SII] 6717, 6731 Å doublet lines are in chip-gaps and are hence unobserved.
cFor these galaxies, the [OII] 3726, 3729 Å doublet lines are in chip-gaps and are hence unobserved.
Note. — Column 1: Name of galaxy; Column 2: galaxy redshift; Column 3-5: Estimated mass, sSFR, and star-formation timescale of galaxies, from their SED fitted to broad-band photometry (see Appendix D); Column 6: Measured Balmer decrement; Column 7: Estimated nebular temperature; Column 8: Estimated electron density; Columns 9–11: Estimated elemental abundances.
aIdentified from broad H lines as AGN host by Harikane et al. (2023).
bFor these galaxies, the abundances were estimated from line fluxes published in the literature.
cFor these galaxies, the ne value has been assumed.
Finally, to further verify the detection of the [ArIII] 7136 Å line fluxes in the publicly available JWST/NIRSPEC spectra for the five highest redshift galaxies in our sample, we computed the stacked 1D spectrum from the individual 1D spectra of CEERS 1536, GLASS 150029, CEERS 1665 and CEERS 1651, adopting the spectral resampling using SpectRes (Carnall 2017) and normalised to their [ArIII] 7136 Å line fluxes. In the rest frame 1D stacked spectrum, the S/N of a detected line should show an increase proportional to the square root of the number of stacked spectra in case of Poissonian noise (e.g. Arnaboldi et al., 2002). The Flux/ in the 1D spectra around the [ArIII] 7136 Å lines, for the four galaxies and their 1D spectral stack is shown in Figure 2b. We find that the [ArIII] 7136 Å line is detected with S/N of 9.27, showing the expected rough increase in S/N compared to the S/N of the line in individual spectra. In the stacked spectrum the He I 7065 Å line is also clearly seen.
Note that for ERO 10612 (), the [ArIII] 7136 Å line lies beyond the red wavelength limit of NIRSPEC, instead both [ArIV] 4711, 4740 Å are independently detected, each with S/N, see Table 1. In the second exposure for ERO 10612 acquired at a different telescope rotator angle, only the [ArIV] 4711 Å line is detected with with S/N. The stacked 1D spectra from the two exposures shows that the Flux/ of the [ArIV] 4711 Å line is higher in the stack (S/N4.68) than in the individual exposures (3.17 and 3.73) by , supporting a robust detection for this line emission.
2.3 Flux measurements and final sample
For the 10 galaxies above with extracted 1D JWST/NIRSPEC grating spectra, flux measurements of the detected emission-lines are then carried out using ALFA (Wesson, 2016). Therefore noise peaks at random wavelengths are never fit, and the wavelengths of all expected fainter lines are known from the bright lines. All lines are fitted to a Gaussian profile iterated during the fit. Since the emission line profile is approximately Gaussian, while the noise peak profile is generally not, even in the unlikely event that a noise peak falls exactly at the wavelength of an expected emission line, a correspondingly large residual will cause a low final S/N, hence a non-detection. The line flux ratios for the emission lines of interest with respect to H are reported in Table 1 for each of the ten galaxies, with 1D spectra of these galaxies shown in Figure 1a.
@@ 2.3 Flux measurements and final sample
In addition to these 10 galaxies, we include three high-redshift SFGs in our sample whose relevant line flux measurements, including ionised O and Ar lines, are published in the literature. One SFG at z1.3 had its O and Ar abundances determined directly based on their [OIII] 4363 Å line detection in NIRSPEC IFU observations (Welch et al., 2024). For two SFGs at z2.2, their direct determination of O abundances were based on Keck/MOSFIRE spectra but with temperature sensitive [OII] 7322,7332 Å line detections, where [ArIII] 7136 Å line fluxes were also reported (Sanders et al., 2023). We use the published line-flux measurements for these three additional SFGs to determine their O & Ar abundances.
In summary, we thus have a sample of 13 emission line galaxies, 3 of them with line flux measurements from the literature, while we measure line fluxes for 10 galaxies from their publicly available 1D JWST/NIRSPEC spectra. For all of them, we determine their O & Ar abundances in this work.